How many liters of antifreeze must be added to 5 liters of water to make no less than a 60% solution and no more than an 80% solution?
Answer by rapaljer(4671)

x = amount of antifreeze at 100%
5 = amount of water at 0%
x+5 = total mixture at 60% or 80%

x = .6(x+5)
x = .6x + 3.0
x - .6x = 3.0
1.0x - .6x = 3.0
.4x = 3.0
x = 3.0/.4 = 30/4 = 7.5 liters for the minimum of 60%

x = .8(x+5)
x = .8x + 4.0
x - .8x = 4.0
1.0x - .8x = 4.0
.2x = 4.0
x = 4.0/.2 = 40/2 = 20 liters for the maximum of 80%

In other words, you should use between 7.5 and 20 liters (inclusive!) of antifreeze.

R^2 at SCC